Use an equation to find the value of x.
There is a scalene triangle that has two sides labeled 7 and 16, and the last side is labeled x. The perimeter is 39.

Answer by ramkikk66(644)

Perimeter = sum of 3 sides = 39
i.e. 7 + 16 + x = 39
x + 23 = 39 -> this is the equation you are looking for.
Solve for x
x = 39 - 23 = 16
As simple as that :)